OpenStudy (anonymous):

Confirm that f(x)=2/x and g(x)=2/x are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (ikram002p):

can u find f(g(x)) and g(f(x)) ?

OpenStudy (ikram002p):

g(x)=2/ f(g(x))=f(2/x) so in f(x) put 2/ instead of x ,got it ?

OpenStudy (ikram002p):

if f(g(x))=g(f(x)) then f(x) and g(x) are inverse to each other

OpenStudy (anonymous):

how would it put it together would it be 2(2/x)/x?..i have the general idea i just don't understand how you plug it into each other

OpenStudy (ikram002p):

will what u did is correct lets simplify it

OpenStudy (ikram002p):

\( \Huge \frac{2}{(\frac{2}{x})}= 2\div(\frac{2}{x}) =2\times(\frac{x}{2}) =x\)

OpenStudy (ikram002p):

so got it nw ?

OpenStudy (mathmale):

In summary: if you use f as the input to g, and the result is x, f and g are inverse functions. You could say the same thing in reverse order: If you use g as the input to f, and if the result is x, f and g are inverse functions.

OpenStudy (anonymous):

 would this be how i would confirm it